# Jin Li    Ye Luo    Xiaowei Zhang

### v1

@@ 5 Conclusion
Our paper examines A/B testing in the presence of spillover effects. Using a multivariate normal framework, we identify conditions that give rise to seesaw experimentation—where continuous adoption of successful innovations leads to performance decline. We show that this phenomenon is more likely to occur when the signal-to-noise ratio is large and correlations across dimensions are more negative. We also show how implementing appropriate hurdle rates can enhance firm performance, and derive optimal thresholds within our analytical framework. At the optimal threshold, the positive gain is equal to the negative externality it imposes.
While our model employs a multivariate normal distribution, the core mechanism we identify extends beyond this specific framework. The seesaw phenomenon and the benefits of hurdle rates persist across various outcome distributions, making our insights broadly applicable. Setting optimal hurdle rates in practice requires careful measurement of innovation outcomes across dimensions, including their correlations. For this purpose, firms can leverage their historical A/B test data, which provides a rich source of information about both direct effects and spillovers.
Our analysis demonstrates that positive hurdle rates emerge as a natural response to negative spillovers across dimensions. This principle extends symmetrically: in environments with positive spillovers, negative hurdle rates would optimally encourage innovation adoption. While the precise value of optimal hurdle rates may vary by context, our findings highlight a fundamental principle: when innovations have multi-dimensional impacts, firms should move beyond zero-threshold A/B testing. The interconnected nature of modern organizations demands a more systematic approach to experimentation—one that explicitly accounts for these cross-dimensional externalities. This insight opens new pathways for firms to design more effective experiments.
Appendix A Proofs for Bivariate Normals
@@
$g(\alpha)\coloneqq\alpha M(\alpha)$ is strictly increasing in $\alpha\in\mathbb{R}$, $g(0)=0$, and $\lim\limits_{\alpha\to\infty}g(\alpha)=1$.
Proof of Proposition 1.
Note: $g(\alpha)=\alpha(1-\Phi(\alpha))/\phi(\alpha)$ and $\phi^{\prime}(\alpha)=(-\alpha)\phi(\alpha)$. Hence,
$g^{\prime}(\alpha)=\frac{1-\Phi(\alpha)+\alpha(-\phi(\alpha))-(-\alpha\phi(\alpha))}{\phi^{2}(\alpha)}=\frac{1-\Phi(\alpha)}{\phi^{2}(\alpha)}>0.$
Moreover, the limit of $g(\alpha)$ can be calculated via L’Hôpital’s rule. ∎
@@
$g(z_{u},z_{v})={}$ $p_{u}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right]$ (3)
@@
The expression (3) can be derived by using the following properties of the truncated normal distribution (Johnson and Kotz 1972, p. 112–113): for any $z$,
@@
By equation (3), a sufficient condition for $g(0,0)<0$ is
$\left\{\begin{array}[]{l}\displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0,\\[10.00002pt] \displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0.\end{array}\right.$ (4)
@@
Appendix B Proofs for Multivariate Normals
@@
which implies $z^{*}$ is the maximizer of $h(z)$. Moreover,

### v3

@@ 5 Conclusion
Our paper examines A/B testing in the presence of spillover effects. Building on a multivariate normal framework, we identify conditions that give rise to seesaw experimentation—where continuous adoption of successful innovations leads to performance decline. We show that this phenomenon is more likely to occur when the signal-to-noise ratio is large and correlations across dimensions are more negative. We also show how implementing appropriate hurdle rates can enhance firm performance, and derive optimal thresholds within our analytical framework. At the optimal threshold, the positive gain is equal to the negative externality it imposes.
While our results are theoretical in nature, they have practical implications: when innovations have multi-dimensional impacts, firms should move away from zero-threshold A/B testing. The interconnected nature of modern organizations demands a more systematic approach to experimentation—one that explicitly accounts for these cross-dimensional externalities. Setting optimal hurdle rates in practice requires careful measurement of innovation outcomes across dimensions, including their correlations. For this purpose, firms can leverage their historical A/B test data, which provides a rich source of information about both direct effects and spillovers.
Appendices
Appendix A Proofs for Bivariate Normal Distribution
@@
$k(\alpha)\coloneqq\alpha M(\alpha)$ is strictly increasing in $\alpha\in\mathbb{R}$, $k(0)=0$, and $\lim\limits_{\alpha\to\infty}k(\alpha)=1$.
Proof of Lemma 1.
Note: $k(\alpha)=\alpha(1-\Phi(\alpha))/\phi(\alpha)$ and $\phi^{\prime}(\alpha)=(-\alpha)\phi(\alpha)$. Hence,
$k^{\prime}(\alpha)=\frac{(1-\Phi(\alpha))(1+\alpha^{2})-\alpha\phi(\alpha)}{\phi(\alpha)}.$
Because $\phi(\alpha)>0$ and $\Phi(\alpha)\in(0,1)$ for all $\alpha$, we have $k^{\prime}(\alpha)>0$ for all $\alpha\leq 0$. Moreover, it is known that the Mills ratio satisfies
$\frac{1-\Phi(\alpha)}{\alpha}>\frac{\alpha}{1+\alpha^{2}},\quad\mbox{for all }\alpha>0;$
see, e.g., Gasull and Utzet, (2014). Hence, $k^{\prime}(\alpha)>0$ for all $\alpha>0$.
Finally, the limit of $k(\alpha)$ can be calculated via L’Hôpital’s rule: ∎
@@
$g(z_{u},z_{v})={}$ $p_{u}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right]$ (A.1)
@@
The expression (A.1) can be derived by using the following properties of the truncated normal distribution (Johnson and Kotz, 1972, p. 112–113): for any $z$,
@@
By equation (A.1), a sufficient condition for $g(0,0)<0$ is
$\left\{\begin{array}[]{l}\displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0,\\[10.00002pt] \displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0.\end{array}\right.$ (A.2)
@@
Appendix B Proofs for Multivariate Normal Distribution
@@
which implies $z^{*}$ is the maximizer of $h(z)$. In addition,
@@
Appendix C Proofs for Bivariate $t$ Distribution
Lemma 4.
Suppose $(U,V)$ follows a symmetric bivariate $t$ distribution. Then,
$f(z)={}$ $2\mu\left(1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)\right)+\sigma(1+\rho)\left(\frac{\delta}{\delta-2}\right)t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right).$
Proof of Lemma 4.
Each marginal ($U$ or $V$) follows a univariate $t$ distribution with mean $\mu$, variance $\sigma^{2}$, and degrees of freedom $\delta>2$. Then, $\mathbb{E}[U|U>z]$ is the expectation of a truncated $t$ distribution with truncation interval $(z,\infty)$. Hence,
$\mathbb{E}[U|U>z]=\mu+\sigma\frac{\delta\Gamma\left(\frac{\delta+1}{2}\right)}{(\delta-1)\sqrt{\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta-1}{2}}\frac{1}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)},$ (C.1)
where $\Gamma$ denotes the gamma function; see, e.g., Ho et al., (2012).
Note that the $t$ density function $t_{\delta}(z;\mu,\sigma^{2})$ is
$t_{\delta}(z;\mu,\sigma^{2})=\frac{\Gamma\left(\frac{\delta+1}{2}\right)}{\sqrt{\sigma^{2}\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta+1}{2}}.$
Therefore,
$t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)={}$ $\frac{\Gamma(\frac{\delta-1}{2})}{\sqrt{\frac{\delta}{\delta-2}(\delta-2)\pi}\Gamma(\frac{\delta-2}{2})}\left(1+\frac{(\frac{z-\mu}{\sigma})^{2}}{\frac{\delta}{\delta-2}(\delta-2)}\right)^{-\frac{\delta-1}{2}}$
[…]
where the last step follows from the gamma function’s property $\Gamma(z+1)=z\Gamma(z)$. Hence, we can rewrite (C.1) as
$\mathbb{E}[U|U>z]=\mu+\sigma\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)}\right).$ (C.2)
The conditional distribution of $V$ given $U$ is a univariate $t$ distribution (see, e.g., Ding, (2016)). In particular, $\mathbb{E}[V|U]=\mu+\rho(U-\mu)$. Thus,
$\mathbb{E}[V|U>z]={}$ $\mathbb{E}\bigl{[}\mathbb{E}[V|U]\big{|}U>z\bigr{]}=\mu+\rho\bigl{[}\mathbb{E}[U|U>z]-\mu\bigr{]}$
[…]
Therefore,
$f(z)={}$ $\mathbb{E}[D(U+V)]$
[…]
where the third equality holds because $\mathbb{E}[U+V|U>z]=\mathbb{E}[U+V|V>z]$ (symmetry), and the last equality follows from (C.2) and (C.3). ∎
Proof of Proposition 7.
The proof follows from Lemma 2 through direct calculations. ∎
Proof of Proposition 8.
Direct calculations based on the expression of $f(z)$ in Lemma 2 yield:
$f^{\prime}(z)={}$ $-\frac{1}{\sigma}t_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)\left[2\mu+(1+\rho)(z-\mu)\right],$
where the following intermediate result is used
[…]
$z^{*}=\frac{\rho-1}{\rho+1}\mu.$
Furthermore, we have
[…]
Thus,
$f^{\prime\prime}(z^{*})=-\frac{1}{\sigma}t_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)(1+\rho)<0,$
proving $z^{*}$ is the maximum of $f(z)$.
Moreover,
$f(z^{*})={}$ $2\mu\left[1-T_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)\right]\left[1+\frac{\sigma(1+\rho)}{2\mu}\left(\frac{\delta}{\delta-2}\right)\frac{t_{\delta-2}\left(\frac{z^{*}-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)}\right]$
[…]
Let $X$ be a standard univariate $t$ distribution. Then, (C.2) implies that
$\mathbb{E}[X|X>\alpha]=\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\alpha;0,1\right)}\right),$
for all $\alpha$. Clearly, $\mathbb{E}[X|X>\alpha]>\alpha$. Hence,
$\frac{1}{\alpha}\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\alpha;0,1\right)}\right)>1.$
for all $\alpha>0$. Setting $(z^{*}-\mu)/\sigma=\alpha$, which is positive by the definition of $z^{*}$, in (C.4) and using the assumption $\mu<0$, we conclude that $f(z^{*})>0$. ∎